Two waves travelling in the same direction are given by: y1 = 0.2 sin⁡(3πx-20t+φ) and y2 = 0.2 sin⁡(3πx-20t), where x and y are in meters and t is in seconds. If the two waves start at the same moment, then the path difference, ∆x, corresponding to a fully destructive interference is:


1/2 m


1/3 m


3/4 m


2/3 m


4/3 m


1
Expert's answer
2021-05-31T15:37:53-0400

"y_1=0.2\sin(3\pi x-20t +\phi)"

"y_2=0.2\sin(3\pi x-20t)"


"y=y_1+y_2"

"y=0.04\sin\bigg(3\pi x-20t+\dfrac{\phi}{2}\bigg)\cos\dfrac{\phi}{2}"

Now,

"k=\dfrac{2\pi}{\lambda}=3\pi"

"\lambda=\dfrac{2\pi}{3\pi}=\dfrac{2}{3}"


Path difference for destructive interference,

"\Delta x=\bigg(n+\dfrac{1}{2}\bigg)\lambda\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space n=0,1,2,3"

for n = 0

"\Delta x=\dfrac{\lambda}{2}=\dfrac{1}{3}"

Therefore path difference, "\Delta x=\dfrac{1}{3}\space m"